OpenStudy (anonymous):

A card is drawn from a standard deck of cards. Find each probability 1. P(jack or queen) 2. P(ace or heart) 3. P(2 or face card)

OpenStudy (anonymous):

1.p=8/52=

OpenStudy (anonymous):

2/13

OpenStudy (anonymous):

2.14/52=7/26

OpenStudy (anonymous):

3.16/52

OpenStudy (anonymous):

1.8/52=2/13 2.16/52 3.16/52

OpenStudy (anonymous):

wow lots of different answers.

OpenStudy (anonymous):

jack or queen. 4 jacks, 4 queens, 8 total.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ace or heart 13 hears and ace of clubs, spades, diamonds. 16 all together

OpenStudy (anonymous):

so 16/52

OpenStudy (anonymous):

4/13

OpenStudy (anonymous):

2 or face card 4 twos, 12 face cards so 16 all together

OpenStudy (anonymous):

so 16/52=4/13

OpenStudy (anonymous):

so we see that my answers weren't even close!

OpenStudy (anonymous):

and mathpro got it on first shot!

OpenStudy (anonymous):

fortunately i can delete my answer so know one knows how wrong i was.